OpenStudy (anonymous):

medal biggie time and fan What is the slope and y-intercept of this line? y = –2x + 8 A. slope: 8, y-intercept: –2 B. slope: 2, y-intercept: –8 C. slope: –1, y-intercept: 4 D. slope: –2, y-intercept: 8

OpenStudy (igreen):

y = mx + b Where m is slope and b is y-intercept.

OpenStudy (igreen):

So what's the slope and y-intercept going to be?

OpenStudy (anonymous):

The formula is y = mx + b. The m is slope, and the b is the y=intercept.

OpenStudy (anonymous):

So what's your answer? @killit!gentlythough.

OpenStudy (anonymous):

I tried doing the math but the fraction got on the wrong side and every thing got mixed up

OpenStudy (igreen):

Um.. you don't have to do any math at all.

OpenStudy (anonymous):

You don't need to do math on it. The slope is m, and the y=intercept is b in the equation y = mx + b.

OpenStudy (igreen):

Just compare it with y = mx + b.

OpenStudy (anonymous):



OpenStudy (anonymous):

There isn't math to do.

OpenStudy (anonymous):

The question uses the equation y = mx + b.

OpenStudy (anonymous):

M is slope, and b is y=intercept. Got it? @killit!gentlythough.

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

wait it's d isn't it?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

yay!